While using a jack a mechanic applied a force of 100N on the effort piston while raising the rear part of a car.Determine the maximum load that can be raised
1
Expert's answer
2019-09-06T09:39:19-0400

We know that

"F=W=mg"

"g=9.8\frac{m}{sec2}"

"F=100\ N"

"m=\frac{F}{g}=\frac{100}{9.8}=10.204\ kg"

So the  the maximum load that can be raised "=10.204\ kg"